OpenStudy (anonymous):

How many 7 digit phone numbers can be formed if the first digit cannot be 0 or 1?

OpenStudy (anonymous):

well, let's write out each posibility for each digit and then multiply

OpenStudy (anonymous):

I was told to use the Counting Principle

OpenStudy (anonymous):

first one can't be 0 or 1 , so we got only 8 possibilities

OpenStudy (anonymous):

8*10*10*10*10*10*10

OpenStudy (anonymous):

the other digits have 10 possibilties so we multiplies them all

OpenStudy (anonymous):

Okay, so we delete 2 outta the 10 possibliity.

OpenStudy (anonymous):

like Ryan did above

OpenStudy (lgbasallote):

you're not planning on stalking someone are you @GOODMAN /:) he

OpenStudy (anonymous):

Wait, how do we know that the others have 10 possibilities, doesnt it apply to the others as well?

OpenStudy (lgbasallote):

jk goodie :DDD

OpenStudy (anonymous):

well, only first digit can't be 0 or 1 ; other digits can be any number from 0 to 9

OpenStudy (anonymous):

Ohhh, okay, got it!

OpenStudy (anonymous):

I need help with one more if you dont mind..

OpenStudy (anonymous):

A photographer is taking a picture o f a bride and groom together with 6 attendants. How many ways can he arrange the 8 people in a line if hte bride and groom stand in the middle?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

little help with photographer problem

OpenStudy (anonymous):

ok ____, ___, ____, B G ____, ____, _____

OpenStudy (anonymous):

\[6!\]

OpenStudy (anonymous):

Thats it?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

that looks right to me

OpenStudy (agreene):

what all of a sudden this has to be in base 10?

OpenStudy (anonymous):

this might surprise you and this might amaze you

OpenStudy (anonymous):

but i am going to knock out joe frazier !

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

btw it is \[6!=6\times 5\times 4\times 3\times 2=720\]

OpenStudy (anonymous):

Yea, i knw what factorials are, but whats is Joe Frazier got to do with this, lol.

OpenStudy (anonymous):

Float like a butterfly, sting like a bee

OpenStudy (anonymous):

His hands can't hit what his eyes can't see.

OpenStudy (lgbasallote):

he beat muhammad ali....

OpenStudy (anonymous):

Hahaha, thats nice.